# Two or three things particle physicists (mis)understand about (pre)heating

### v1

@@ 2 The poor particle-physics approach
The usual Boltzmann framework for studying (re)heating focuses on tracking the phase-space distribution of particles produced by perturbative decays or scatterings of the inflaton condensate by solving sets of BEQs [7]. This approach enables a systematic analysis of the time evolution of the energy densities associated with the inflaton, radiation, and any other exotic states generated during (re)heating. The Boltzmann framework has been extensively discussed in the literature, particularly within the particle physics community; see, e.g., Refs. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Here, we provide a concise overview, emphasizing the key assumptions underlying this intrinsically perturbative treatment.
Immediately after inflation, the inflaton field $\phi$ is typically spatially homogeneous, with any residual inhomogeneities or self-interactions considered negligible. The post-inflationary evolution of the Universe is then primarily dictated by the inflaton potential, with minimal contributions from other energy components, until radiation domination takes over. Among the various inflationary models consistent with observations [37], ranging from the seminal Starobinsky model [38, 39, 40, 41] to Higgs inflation [42] and its variants [43, 44, 45, 46, 47] or other universal attractors [48, 49], we focus here for concreteness on a $T$-model of inflation with potential [48]
@@
with the dots indicating derivatives with respect to coordinate time $t$, and the commas derivatives with respect to the field $\phi$. Here, $H$ represents the Hubble expansion rate, while $\Gamma_{\phi}$ corresponds to the total inflaton decay rate, which quantifies the energy transfer rate from individual inflaton quanta to the conventional matter and radiation components. This decay rate enters the equation of motion as an additional friction-like term, effectively governing the dissipation of the inflaton energy density over time. In most particle physics approaches, this decay rate is assumed to be either constant or only weakly dependent on time, allowing for a gradual and controlled energy transfer that smoothly connects inflationary (re)heating with the subsequent thermal history of the Universe. The exact form of the decay rate depends on the particular inflaton decay mode, which, in turn, is influenced by the underlying particle-physics model.
@@
In the context of (re)heating, the interactions of the inflaton with both SM and BSM fields can be broadly categorized into two primary mechanisms: direct decay and scatterings. On the one hand, in the decay channel, the inflaton field $\phi$ can transition into pairs of lighter particles via trilinear interactions, such as $\mu\,\phi\,|\varphi|^{2}$ for scalar fields $\varphi$ (e.g., the Higgs boson doublet) or Yukawa-like couplings $y_{\psi}\,\phi\,\overline{\Psi}\,\Psi$ for fermionic fields $\Psi$. These interactions lead to two-body decays, with the coupling parameters $\mu$ and $y_{\psi}$ controlling their respective strengths. On the other hand, scattering processes, which provide an alternative pathway for energy transfer, occur via quartic couplings of the form $g\,\phi^{2}\,|\varphi|^{2}$, allowing 2-to-2 interactions of strength $g$ that populate the thermal bath with SM and BSM particles. Beyond this model-dependent channel, there exists an irreducible gravity-mediated production following from the unavoidable interaction term $\sqrt{-g}\,\mathcal{L}_{\rm grav}\supset-2/M_{P}\,h_{\mu\nu}\,T^{\mu\nu}$ between the graviton field $h_{\mu\nu}$ and the energy-momentum tensor $T_{\mu\nu}$ of all matter fields in the theory [51], with $M_{P}=(8\pi\,G_{N})^{-1/2}\simeq 2.4\times 10^{18}$ GeV the reduced Planck mass. In this case, (re)heating can occur through 2-to-2 scatterings of the inflaton condensate into SM final states [52, 53, 54, 14, 55, 21, 25, 27, 56, 57], provided of course that the available energy density is large enough to partially overcome the involved Planck suppression. Alternatively, gravitational (re)heating can also occur through gravity-mediated decays as in the case of linear non-minimal interactions of the inflaton [58].
@@
Given the above set of interactions and assuming that the bare masses of the decay products are significantly smaller than that of the inflaton condensate, the inflaton’s total decay width, computed using standard Feynman rules techniques, can be approximated as [59, 60]
$\Gamma_{\phi}=\frac{\gamma^{2}_{\text{eff}}}{8\pi}\times\begin{dcases}m_{\phi}(a)&\quad\text{for fermionic decay},\\ m^{-1}_{\phi}(a)&\quad\text{for bosonic decay},\\ \rho_{\phi}/m_{\phi}^{3}&\quad\text{for bosonic scattering},\end{dcases}$ (2.3)
for explicit couplings to matter, and as [24, 27, 56]
@@
for minimal gravitational couplings, with $\alpha_{n}$ encoding the effect of oscillating inflaton condensate. Here, $\rho_{\phi}$ is the inflaton energy density,
@@
stands for the inflaton mass in the vicinity of the origin ($|\phi|\ll\Lambda$), and
${\gamma^{2}_{\text{eff}}}=\gamma^{2}\,(n-1)\,(n+2)\left(\frac{\omega}{m_{\phi}}\right)^{q}\,\sum_{j=1}^{\infty}j^{q}\,|\mathcal{P}j|^{2}\,\left\langle\left[1-\left(\frac{2\,m_{\phi}}{j\,\omega}\right)^{2}\mathcal{P}\right]^{q/2}\right\rangle$ (2.6)
@@
where $\mathcal{P}(t)$ is a quasi-periodic, fast-oscillating function, and $\phi_{0}(t)$ is a slowly-varying envelope. Note that for $n=2$, both the inflaton mass and the frequency $\omega$ in this mode decomposition are exactly constant, while displaying an explicit field dependence for $n>2$ which makes them decrease with time.
@@
where $p$ collectively denotes all the final particles produced from inflaton decay or scattering, $d\Pi$’s are the Lorentz-invariant phase space, |¯$|\overline{\mathcal{M}}|^{2}$ is the spin-averaged squared amplitude for the underlying process concerned (with any symmetry factors included), and $\left(1\pm f_{p}\right)$ takes care of Pauli-blocking/stimulated emission effects for the product particles. Clearly, all information from the underlying particle physics model is imprinted in the matrix element $\mathcal{M}$, which is computed perturbatively for a given decay/scattering diagram involving inflaton and SM fields. Now, assuming $i)$ no backreaction producing inflaton condensate from the bath, i.e., |¯$|\overline{\mathcal{M}}|^{2}_{p\to\phi}=0$, and $ii)$ that the Bose enhancement and Pauli blocking effects due to the decay products are negligible (that is, $1\pm f_{p}\simeq 1$), one can arrive at the evolution equation for the inflaton energy density $\rho_{\phi}$, which reads [65]
@@
with $\alpha$ depending on the inflaton potential and the details fo the energy transfer (spin of the daughter particles and annihilation/decay channel). In any case, to ensure the success of BBN, the (re)heating temperature must satisfy $T_{\text{rh}}>T_{\text{BBN}}\simeq 4$ MeV [72, 73, 74, 75, 76, 77].
@@ 3 Here be dragons
Almost three decades ago, Kofman, Linde, and Starobinsky [110, 111] (see also Refs. [112, 113, 64, 114]) identified fundamental shortcomings in the widely used Boltzmann approach presented in the previous section, demonstrating that it is not just inaccurate, but also fundamentally inadequate, at least during the initial stages of (re)heating; cf. Fig. 3. First, Eq. (2.2) violates the fluctuation-dissipation theorem, which states that dissipation within a system necessarily generates fluctuations [115, 116]. In particular, the analysis does not account for how these fluctuations modify the effective mass of the inflaton condensate. Another issue is that the perturbative approximation employed when computing amplitudes and decay rates neglects both the coherent nature of the inflaton field and crucial phenomena such as Bose enhancement and Pauli blocking effects. Most critically, even for coupling values where radiative corrections still remain negligible, perturbative methods break down. For example, if the inflaton oscillates in a potential steeper than quadratic, homogeneous oscillations can trigger explosive instabilities leading to rapid amplification of field fluctuations. Unlike the slow, steady energy leakage of perturbative decays, this mechanism violently disrupts the inflaton condensate, causing its fragmentation and initiating a highly non-linear interplay with its decay products, which can also undergo non-perturbative enhancements, further complicating the dynamics and often requiring the use of classical lattice simulations for a proper treatment; see e.g. Refs. [117, 118, 119]. [^3] For a justification of the validity of the classical analysis and its limitations, see e.g. Refs. [120, 121, 122]. This section explores these dramatic, non-perturbative phenomena that challenge the conventional perturbative picture presented in Section 2.
@@ 3.1 Inflaton self-resonance and fragmentation
At the end of inflation, the inflaton field undergoes coherent oscillations around the minimum of its potential. In models where the potential is of the form (2.1) with $n>2$, these oscillations can trigger self-resonance instabilities, driving the energy transfer from the homogeneous condensate $\bar{\phi}(t)$ into spatial gradients and ultimately leading to its fragmentation. This instability arises because the oscillating field sources its own perturbations $\delta\phi(t,\vec{x})=\phi(t,\vec{x})-\bar{\phi}(t)$ amplifying them non-perturbatively [123, 124, 125, 126, 127, 128]. In Fourier space, we have
@@
Although the (pre)heating dynamics is often thought to be primarily dictated by the shape of the potential near its minimum, surprising effects can emerge when the potential becomes shallower than quadratic at large field values. In such cases, fragmentation can give rise to localized, long-lived structures known as oscillons [130, 131, 132]: pseudostable, non-topological solitons that emerge from the nonlinear dynamics of the inflaton field, cf. Fig. 4. For instance, while the condensate does not fragment in $T$-models of inflation with $n=2$ and $\Lambda\sim M_{P}$, [^4] Note that this implicitly disregards gravitational effects, which have been shown to ultimately lead to the formation of nonlinear structures, even in free field cases, much like the gravitational instabilities seen in pressureless matter in the late Universe [133, 134]. long-lived oscillon configurations form efficiently for $\Lambda\ll M_{P}$, collectively acting as dust and dominating the energy background for many $e$-folds of post-inflationary expansion [135] (see also Refs. [136, 137]). However, for potentials steeper than quadratic ($n>2$), oscillons are generally not formed, and fragmentation leads to transient structures that decay more rapidly [135, 138], giving rise to a radiation-dominated Universe.
@@ 3.3 Multifield blocking effects
The picture presented in the previous section may change significantly in the presence of multiple scalar fields $\phi_{I}$ that contribute to inflation. In such multifield setups, each inflaton component typically oscillates with a different mass and amplitude, leading to a superposition of oscillatory modes. If all the fields are coupled to a common matter species—for instance, through interaction terms like $\sum_{I}g_{I}^{2}\,\phi_{I}^{2}\,\varphi^{2}$—the resulting effective mass of the daughter field $\varphi$ becomes a time-dependent function with contributions from all $\phi_{I}(t)$ components. Due to the differing frequencies, phases, and decay rates of these inflaton oscillations, the combined evolution often lacks the regularity required for efficient resonance. This loss of coherence between the inflaton fields—sometimes referred to as de-phasing—can “block” the amplification of the daughter field’s fluctuations by smoothing out the zero crossings of its effective mass. As a result, sharp adiabaticity violations necessary for triggering explosive particle production may be suppressed. This phenomenon reduces the efficiency of (pre)heating, delaying energy transfer to the radiation bath, and potentially prolonging the (re)heating phase.
Despite the aforementioned suppression, multifield models are not necessarily insensitive to non-perturbative processes. The structure of the couplings plays a critical role. For example, introducing trilinear couplings or allowing negative couplings for certain inflaton components can restore or even enhance the resonant instabilities, bypassing the blocking effects associated with simple quadratic couplings.

### v2

@@ 2 The poor particle-physics approach
The usual Boltzmann framework for studying (re)heating focuses on tracking the phase-space distribution of particles produced by perturbative decays or scatterings of the inflaton condensate by solving sets of BEQs [7]. This approach enables a systematic analysis of the time evolution of the energy densities associated with the inflaton, radiation, and any other exotic states generated during (re)heating. The Boltzmann framework has been extensively discussed in the literature, particularly within the particle physics community; see, e.g., Refs. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Here, we provide a concise overview, emphasizing the key assumptions underlying this intrinsically perturbative treatment.
As is customary, we assume that the inflaton’s self-interactions and its couplings to both the SM and the BSM fields are sufficiently small so as not to significantly alter the inflationary dynamics prior to the end of inflation. This cold-inflation assumption sets our analysis apart from warm-inflation scenarios [38, 39, 40, 41, 42, 43], where dissipative effects play a central role throughout the inflationary period. Within this framework, any decay products generated during inflation are exponentially diluted by the rapid expansion of the Universe. Consequently, the inflaton field remains effectively homogeneous at the end of inflation, with residual inhomogeneities considered negligible. The post-inflationary evolution of the Universe is then primarily dictated by the inflaton potential, with minimal contributions from other energy components, until radiation domination takes over. Among the various inflationary models consistent with observations [44], ranging from the seminal Starobinsky model [45, 46, 47, 48] to Higgs inflation [49] and its variants [50, 51, 52, 53, 54] or other universal attractors [55, 56], we focus here for concreteness on a $T$-model of inflation with potential [55]
@@
with the dots indicating derivatives with respect to coordinate time $t$, and the commas derivatives with respect to the field $\phi$. Here, $H$ represents the Hubble expansion rate, while $\Gamma_{\phi}$ corresponds to the total inflaton decay rate, which quantifies the energy transfer rate from individual inflaton quanta to the conventional matter and radiation components. This decay rate enters the equation of motion as a local friction-like term, effectively governing the dissipation of the inflaton energy density over time. In most particle physics approaches, this decay rate is assumed to be either constant or only weakly dependent on time, allowing for a gradual and controlled energy transfer that smoothly connects inflationary (re)heating with the subsequent thermal history of the Universe. The exact form of the decay rate depends on the particular inflaton decay mode, which, in turn, is influenced by the underlying particle-physics model.
@@
In the context of (re)heating, the interactions of the inflaton with both SM and BSM fields can be broadly categorized into two primary mechanisms: direct decay and scatterings. On the one hand, in the decay channel, the inflaton field $\phi$ can transition, for instance, into pairs of lighter particles via trilinear interactions, such as $\mu\,\phi\,|\varphi|^{2}$ for scalar fields $\varphi$ (e.g., the Higgs boson doublet) or Yukawa-like couplings $y_{\psi}\,\phi\,\overline{\Psi}\,\Psi$ for Dirac fermionic fields $\Psi$. These interactions lead to two-body decays, with the coupling parameters $\mu$ and $y_{\psi}$ controlling their respective strengths. Depending on the underlying particle physics model, the main decay channel for the inflaton can alternatively involve other spin states or additional three or even four bodies, as discussed, for example, in Refs. [58, 59]. On the other hand, scattering processes, which provide an alternative pathway for energy transfer occur, for instance, via quartic couplings of the form $g\,\phi^{2}\,|\varphi|^{2}$, allowing 2-to-2 interactions of strength $g$ that populate the thermal bath with SM and BSM particles. Beyond this model-dependent channel, there exists an irreducible gravity-mediated production following from the unavoidable interaction term $\sqrt{-g}\,\mathcal{L}_{\rm grav}\supset-2/M_{P}\,h_{\mu\nu}\,T^{\mu\nu}$ between the graviton field $h_{\mu\nu}$ and the energy-momentum tensor $T_{\mu\nu}$ of all matter fields in the theory [60], with $M_{P}=(8\pi\,G_{N})^{-1/2}\simeq 2.4\times 10^{18}$ GeV the reduced Planck mass. In this case, (re)heating can occur through 2-to-2 scatterings of the inflaton condensate into SM final states [61, 62, 63, 14, 64, 21, 25, 27, 65, 66], provided of course that the available energy density is large enough to partially overcome the involved Planck suppression. Alternatively, gravitational (re)heating can also occur through gravity-mediated decays as in the case of linear non-minimal interactions of the inflaton [67].
@@
Given the above set of interactions and assuming that the effective masses of the decay products are smaller than that of the inflaton condensate, [^2] Note in particular that this implicitly disregards potentially large time-dependent masses for the product particles induced by the inflaton field expectation value, as will be further emphasized in Section 3. the inflaton’s total decay width, computed using standard Feynman rules techniques, can be approximated as [68, 69]
$\Gamma_{\phi}=\frac{\gamma^{2}_{\text{eff}}}{8\pi}\times\begin{dcases}m_{\phi}(a)&\quad\text{for decays into Dirac fermions},\\ m^{-1}_{\phi}(a)&\quad\text{for decays into complex doublet scalars},\\ \rho_{\phi}/m_{\phi}^{3}&\quad\text{for annihilations into complex doublet scalars},\end{dcases}$ (2.3)
for explicit couplings to matter. In the case of minimal gravitational couplings [24, 27, 65]
@@
with $\alpha_{n}$ encoding the effect of oscillating inflaton condensate. Here, $\rho_{\phi}$ is the inflaton energy density,
@@
stands for the inflaton mass in the vicinity of the origin ($|\phi|\ll\Lambda$),
${\gamma^{2}_{\text{eff}}}=\gamma^{2}\,(n-1)\,(n+2)\left(\frac{\omega}{m_{\phi}}\right)^{q}\,\sum_{j=1}^{\infty}j^{q}\,|\mathcal{P}j|^{2}\,\left\langle\left[1-\left(\frac{2\,m_{p}}{j\,\omega}\right)^{2}\mathcal{P}\right]^{q/2}\right\rangle$ (2.6)
@@
where $\mathcal{P}(t)$ is a quasi-periodic, fast-oscillating function, and $\phi_{0}(t)$ is a slowly-varying envelope. Note that in Eq. (2.6) the factor in the edgy brackets correspond to an average over the kinematical factors. For $n=2$, both the inflaton mass and the frequency $\omega$ in this mode decomposition are exactly constant, while displaying an explicit field dependence for $n>2$ which makes them decrease with time.
@@
where $p$ collectively denotes all the final particles produced from inflaton decay or scattering, $d\Pi$’s are the Lorentz-invariant phase space, |¯$|\overline{\mathcal{M}}|^{2}$ is the spin-averaged squared amplitude for the underlying process concerned (with any symmetry factors included), and $\left(1\pm f_{p}\right)$ takes care of Pauli-blocking/stimulated emission effects for the product particles. Clearly, all information from the underlying particle physics model is imprinted in the matrix element $\mathcal{M}$, which is computed perturbatively for a given decay/scattering diagram involving inflaton and SM fields. Now, assuming $i)$ no production of inflaton condensate from the bath, i.e., |¯$|\overline{\mathcal{M}}|^{2}_{p\to\phi}=0$, and $ii)$ that the Bose enhancement and Pauli blocking effects due to the decay products are negligible (that is, $1\pm f_{p}\simeq 1$), one can arrive at the evolution equation for the inflaton energy density $\rho_{\phi}$, which reads [74]
@@
with $\alpha$ depending on the inflaton potential and the details fo the energy transfer (spin of the daughter particles and annihilation/decay channel). In any case, to ensure the success of BBN, the (re)heating temperature must satisfy $T_{\text{rh}}>T_{\text{BBN}}\simeq 4$ MeV [81, 82, 83, 84, 85, 86, 87].
@@ 3 Here be dragons
Almost three decades ago, Kofman, Linde, and Starobinsky [120, 121] (see also Refs. [122, 123, 73, 124]) identified fundamental shortcomings in the widely used Boltzmann approach presented in the previous section, demonstrating that it is not just inaccurate, but also fundamentally inadequate, at least during the initial stages of (re)heating; cf. Fig. 3. First, Eq. (2.2) violates the fluctuation-dissipation theorem, which states that dissipation within a system necessarily generates fluctuations [125, 126]. In particular, the analysis does not account for how these fluctuations modify the effective mass of the inflaton condensate. Furthermore, the use of a local decay term relies on the adiabatic-Markovian approximation, which is valid when the timescales associated with the decay process are much shorter than the Hubble time, which might not apply at at the onset of reheating or more generically in warm inflationary scenarios (see e.g. Refs. [127, 128]). Another issue is that the perturbative approximation employed when computing amplitudes and decay rates neglects both the coherent nature of the inflaton field and crucial phenomena such as Bose enhancement and Pauli blocking effects. Most critically, even for coupling values where radiative corrections still remain negligible, perturbative methods break down. For example, if the inflaton oscillates in a potential steeper than quadratic, homogeneous oscillations can trigger explosive instabilities leading to rapid amplification of field fluctuations. Unlike the slow, steady energy leakage of perturbative decays, this mechanism violently disrupts the inflaton condensate, causing its fragmentation and initiating a highly non-linear interplay with its decay products, which can also undergo non-perturbative enhancements, further complicating the dynamics and often requiring the use of classical lattice simulations for a proper treatment; see e.g. Refs. [129, 130, 131]. [^4] For a justification of the validity of the classical analysis and its limitations, see e.g. Refs. [132, 133, 134]. This section explores these dramatic, non-perturbative phenomena that challenge the conventional perturbative picture presented in Section 2.
@@ 3.1 Inflaton self-resonance and fragmentation
At the end of inflation, the inflaton field undergoes coherent oscillations around the minimum of its potential. In models where the potential is of the form of Eq. (2.1) with $n>2$, these oscillations can trigger self-resonance instabilities, driving the energy transfer from the homogeneous condensate $\bar{\phi}(t)$ into spatial gradients and ultimately leading to its fragmentation. [^5] Note that this implicitly disregards gravitational effects, which have been shown to ultimately lead to the formation of nonlinear structures, even in free field cases ($n=2$), much like the gravitational instabilities seen in pressureless matter in the late Universe [135, 136]. This instability arises because the oscillating field sources its own perturbations $\delta\phi(t,\vec{x})=\phi(t,\vec{x})-\bar{\phi}(t)$ amplifying them non-perturbatively [137, 138, 139, 140, 141, 142]. In Fourier space, we have
@@
Although the (pre)heating dynamics is often thought to be primarily dictated by the shape of the potential near its minimum, surprising effects can emerge when the potential becomes shallower than quadratic at large field values. In such cases, fragmentation can give rise to localized, long-lived structures known as oscillons [144, 145, 146]: pseudostable, non-topological solitons that emerge from the nonlinear dynamics of the inflaton field, cf. Fig. 4. For instance, while the condensate does not fragment in $T$-models of inflation with $n=2$ and $\Lambda\sim M_{P}$, [^6] We refer again the reader to footnote 5. long-lived oscillon configurations form efficiently for $\Lambda\ll M_{P}$, collectively acting as dust and dominating the energy background for many $e$-folds of post-inflationary expansion [147] (see also Refs. [148, 149]). However, for potentials steeper than quadratic ($n>2$), oscillons are generally not formed, and fragmentation leads to transient structures that decay more rapidly [147, 150], giving rise to a radiation-dominated Universe.
@@ 3.3 Multifield blocking effects
The picture presented in the previous section may change significantly in the presence of multiple scalar fields $\phi_{I}$ that contribute to inflation. In such multifield setups, each inflaton component typically oscillates with a different mass and amplitude, leading to a superposition of oscillatory modes. If all the fields are coupled to a common matter species—for instance, through interaction terms like $\sum_{I}g_{I}^{2}\,\phi_{I}^{2}\,\varphi^{2}$—the resulting effective mass of the daughter field $\varphi$ becomes a time-dependent function with contributions from all $\phi_{I}(t)$ components. Due to the differing frequencies, phases, and decay rates of these inflaton oscillations, the combined evolution often lacks the regularity required for efficient resonance. This loss of coherence between the inflaton fields—sometimes referred to as de-phasing—can “block” the amplification of the daughter field’s fluctuations by smoothing out the zero crossings of its effective mass [161, 162]. As a result, sharp adiabaticity violations necessary for triggering explosive particle production may be suppressed. This phenomenon reduces the efficiency of (pre)heating, delaying energy transfer to the radiation bath, and potentially prolonging the (re)heating phase.
Despite the aforementioned suppression, multifield models are not necessarily insensitive to non-perturbative processes. The structure of the couplings plays a critical role. For example, introducing trilinear couplings or allowing negative couplings for certain inflaton components can restore or even enhance the resonant instabilities, bypassing the blocking effects associated with simple quadratic couplings [155].